OpenStudy (anonymous):

Lashonda found a recipe that makes 2 1/2 dozen cupcakes and requires 1/4 cup of vegetable oil. If she makes 90 cupcakes for a family reunion picnic, how much vegetable oil will she need?

OpenStudy (jack1):

"Lashonda found a recipe that makes 2 1/2 dozen cupcakes" ...2 half dozen cupcakes... so what you're really saying is a dozen cupcakes... yeah?

OpenStudy (anonymous):

Yeah

OpenStudy (jack1):

ok, then for every 12 cupcaakes u make, u will need 0.25 cups of veggie oil so as a ratio, that's: \[ \huge \frac{\frac{1}{4} cups}{12 cakes} \] so if you want 90 cakes, multiply it bu our ratio to get your answer in oil \[ \huge \frac{\frac{1}{4} cups}{12 cakes} \times \frac {90 cakes}{1}\] cakes cancel out top and bottom, so your left with just oil multiply across \[ \huge \frac{\frac{1}{4} cups \times 90 }{12 \times 1} \] \[ \huge \frac{\frac{90}{4} cups }{12 } \] \[ \huge \frac{\frac{45}{2} cups }{12 } \] \[ \huge \frac{15}{8 } cups ...approx= 1.88 cups -ish\]

OpenStudy (jack1):

i see the choice...

OpenStudy (jack1):

a) 1/4 of a cup will make...12 cakes

OpenStudy (jack1):

b) 1/2 cup will make ... 24 cakes

OpenStudy (jack1):

c) 3/4 cups will make ...36 cakes... so:

OpenStudy (jack1):

so does that all make sense?